OpenStudy (anonymous):

The equation of line CD is y = -2x - 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5). y = -2x + 13 y = x + 7 y = x + 3 y = − 2x − 3

OpenStudy (butterflydreamer):

Firstly can you identify the slope of line CD? [remember: y = mx + b where m = slope] So for y = -2x - 2 , what would m/slope be ?

OpenStudy (anonymous):

-2x?

OpenStudy (butterflydreamer):

you're close but the "x" part won't be included for your slope :) It's just -2. Like in your general formula : y = mx + b . ONLY the "m" is your slope. So for y = -2x -2 , m = -2 :)

OpenStudy (anonymous):

u must resolve it with point x = 4, y =5 y=x+7 5 = 4+7 5=11 its not correct

OpenStudy (butterflydreamer):

Okay so now since we know that the slope of CD " -2 " We want to write an equation of a line that is parallel to CD and contains point (4,5) Recall that the slope-intercept form is this :